# Channel Dependence, Limited Lookback Windows, and the Simplicity of Datasets: How Biased is Time Series Forecasting?

### v2

@@ Appendix G Full Lookback Window Analysis
Figure 4. Analysis of Frequencies of Best Performing Lookback Windows Grouped by : (a) datasets, (b) CI models, and (c) CD models. All recorded for the forecasting horizon 96.
@@
Figure 5. Analysis of Frequencies of Best Performing Lookback Windows Grouped by : (a) datasets, (b) CI models, and (c) CD models. All recorded for the forecasting horizon 192.
@@
Figure 6. Analysis of Frequencies of Best Performing Lookback Windows Grouped by : (a) datasets, (b) CI models, and (c) CD models. All recorded for the forecasting horizon 336.
@@
Figure 7. Analysis of Frequencies of Best Performing Lookback Windows Grouped by : (a) datasets, (b) CI models, and (c) CD models. All recorded for the forecasting horizon 720.
Appendix H ODE Datasets Sample Visualization
In this section, a visualization is shown for the rest of the 6 ODE datasets used in this research. In Figure˜8, a visualization of a 1000 timesteps from cell cycle based ODE is shown. Two different cyclin degradation activation are shown with channel 2 (middle top) activating the degradation of cyclin C1 levels (middle bottom). Same phenomena is shown clearly on the right 2 channels of the figure for cyclin C2 levels. In Figure˜9Figure˜10Figure˜11, and Figure˜12, we can see a visualization of the first 1000 time steps of the blinking rotlet, hopfield, lorenz and lorenz coupled time series based datasets. We can see clearly on these figures, how the dynamics of these different datasets generate complex yet real-world like time series data, with highs and lows from one channel triggering highs and lows in other channels.
@@ Appendix H ODE Datasets Sample Visualization
Figure 8. Time series visualization of the cell cycle ODE-based time series sample.
@@
Figure 9. Time series visualization of the blinking rotlet ODE-based time series sample.
Figure 10. Time series visualization of the hopfield ODE-based time series sample.
Figure 11. Time series visualization of the lorenz ODE-based time series sample.
@@
Figure 12. Time series visualization of the Lorenz Coupled ODE-based time series sample.
Appendix I Cell Cycle ODE Discussion
We discussed already the interesting properties of Double Pendulum ODEs. In this part of the appendix, we would like to extend more the discussion on the coupling between the ODEs and how variables interact across time. Equations 1-6 presents the system of ODEs, we used to generate the cell cycle dataset (Romond et al., 1999). These 6 equations show two different oscillators based on the concentration levels of cyclin ($C_{1}$ and $C_{2}$) and how they motivate the progression of the cell cycle. The cyclin concentration reflects directly on the activated Cyclin Dependent Kinases (cdk1, cdk2) which represented in the equations by $M_{1}$ and $M_{2}$ respectively. It also affect directly enzymes $X_{1}$ and $X_{2}$. $v_{ij}$ and $v_{dj}$ are constants that are multiplied in equation 1 for $j=1$ and equation 4 for $j=2$ as a rate of cyclin synthesis. In equations 1, we can see clearly the coupling between the ODE variables $C_{1}$ and $M_{2}$. Whereas, in equation 4, we can see the coupling between $C_{2}$ and $M_{1}$. All of this shows how information from one channel can cause spikes or change of trend in other channels, showing clear channel dependence in the generated time series data. Also, this section is valuable to show the very important application of such equation on a real life biomedical process which elevates the importance of our study more. Please note that these equations were taken from (Romond et al., 1999) and for more details, we encourage the reader to go through their work as well.
(8) $\frac{dC_{1}}{dt}=\frac{v_{i1}K_{im1}}{K_{im1}+M_{2}}-\frac{v_{d1}X_{1}C_{1}}{K_{d1}+C_{1}}-k_{d1}C_{1}$
(9) $\frac{dM_{1}}{dt}=\frac{V_{1}\left(1-M_{1}\right)}{K_{1}+\left(1-M_{1}\right)}-\frac{V_{2}M_{1}}{K_{2}+M_{1}}$
(10) $\frac{dX_{1}}{dt}=\frac{V_{3}\left(1-X_{1}\right)}{K_{3}+\left(1-X_{1}\right)}-\frac{V_{4}X_{1}}{K_{4}+X_{1}}$
(11) $\frac{dC_{2}}{dt}=\frac{v_{i2}K_{im2}}{K_{im2}+M_{1}}-\frac{v_{d2}X_{2}C_{2}}{K_{d2}+C_{2}}-k_{d2}C_{2}$
(12) $\frac{dM_{2}}{dt}=\frac{U_{1}\left(1-M_{2}\right)}{H_{1}+\left(1-M_{2}\right)}-\frac{U_{2}M_{2}}{H_{2}+M_{2}}$
(13) $\frac{dX_{2}}{dt}=\frac{U_{3}\left(1-X_{2}\right)}{H_{3}+\left(1-X_{2}\right)}-\frac{U_{4}X_{2}}{H_{4}+X_{2}}$
(14) $V_{1}=\frac{C_{1}}{K_{c1}+C_{1}}V_{M_{1}},\qquad V_{3}=M_{1}\cdot V_{M_{3}}$
(15) $U_{1}=\frac{C_{2}}{K_{c2}+C_{2}}U_{M_{1}},\qquad U_{3}=M_{2}\cdot UM_{3}$
Appendix J Ablation Study on RevIN
@@ Appendix I Cell Cycle ODE Discussion
In this section, we include an important ablation study where we test the importance of using Reversible Instance Normalization (RevIN)  (Kim et al., 2021) on 2 standard datasets and 2 datasets from the chaotic benchmark. Crossformer was originally proposed without RevIN, so we added the RevIN component to both the CI and CD versions to measure how much of an effect window size this can have on different types of datasets. Moreover we included one more strong baseline in PatchTST in the study measuring the effect of RevIN on its main CI version.
We got 2 interesting patterns that can be seen also from table  Table˜15 which are the following:
@@
  • Introducing revin helps to improve the performance on the both variants of Crossformer on the standard datasets.
  • On the contrary on the chaotic ODEs the model without revin is performing better than its counterpart with revin.
@@
These findings shows that in general as confirmed in the literature RevIN can help in the performance of different model on standard datasets. While this being not totally true for PatchTST, it has been true for most models in the literature  (Chen et al., 2023). Important as emphasized throughout this paper, that is not easy task to generalize to different types of datasets, and that is no different for RevIN as it deteriorates performance if introduced on Crossformer on the chaotic ODE datasets. This shows that the different nature of Chaotic ODEs with the complex patterns can not benefit from the normalization+denormalization technique introduced in RevIN.
Dataset / Model PatchTST_CI CrossFormer_CD CrossFormer_CI
with RevIN w/o RevIN with RevIN w/o RevIN with RevIN w/o RevIN
@@
Table 15. This table shows an ablation study on the reversible instance normalization technique  (Kim et al., 2021) which showed to improve performance for most model on the standard datasets. The best variant of model is in bold and best overall for each dataset is highlighted in blue. These results are based on the forecasting horizon 96.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
PatchTST CD-96 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
@@
PatchTST CI-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
PatchTST CD-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
TSMixer CI-192 lr:0.0005,num_blocks:8,hidden_size:1.02e+03,dropout:0.705,activation:relu,seq_len:512
@@
PatchTST CI-336 lr’:1.96e-07, ’e_layers’:5, ’d_ff’: 256, ’d_model’:1024, ’dropout’:0.295,’fc_dropout’:0.622, ’patch_size’:8,’stride’:8,’seq_len’:720
PatchTST CD-336 lr’: 9.1666e-05, ’e_layers’: 6, ’d_ff’: 1024, ’d_model’: 1024, ’dropout’: 0.086, ’fc_dropout’: 0.5099, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
@@
PatchTST CI-720 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
@@
TSMixer CI-720 lr:0.0004,num_blocks:8,hidden_size:1.02e+03,dropout:0.754,activation:relu,seq_len:512
@@
Table 16. Hyperparameter settings for the BlinkingRotlet dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:0.0003,e_layers:10,d_ff:256,d_model:128,dropout:0.081,fc_dropout:0.390,patch_size:8,stride:8,seq_len:336
@@
PatchTST CI-336 lr’: 0.00027, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 128, ’dropout’: 0.081, ’fc_dropout’: 0.390, ’patch_size’: 8, ’stride’: 8, ’seq_len’: 336
PatchTST CD-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.53896, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
@@
PatchTST CI-720 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
PatchTST CD-720 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
@@
Table 17. Hyperparameter settings for the CellCycle dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
@@
PatchTST CI-192 lr:1.53e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.112,fc_dropout:0.457,patch_size:16,stride:4,seq_len:192
PatchTST CD-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
@@
TSMixer CD-192 lr:0.0032,num_blocks:7,hidden_size:1.02e+03,dropout:0.006,activation:relu,seq_len:192
@@
PatchTST CI-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.539, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
PatchTST CD-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.5389, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
TSMixer CI-336 lr:0.0012,num_blocks:8,hidden_size:1.02e+03,dropout:0.307,activation:relu,seq_len:192
@@
PatchTST CD-720 lr:0.0002,e_layers:1,d_ff:1.02e+03,d_model:512,dropout:0.471,fc_dropout:0.710,patch_size:8,stride:8,seq_len:336
TSMixer CI-720 lr:0.0005,num_blocks:7,hidden_size:1.02e+03,dropout:0.169,activation:relu,seq_len:192
@@
Table 18. Hyperparameter settings for the DoublePendulum dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:0.0001,e_layers:1,d_ff:512,d_model:256,dropout:0.572,fc_dropout:0.631,patch_size:16,stride:4,seq_len:96
@@
PatchTST CI-192 lr:7.27e-06,e_layers:4,d_ff:1.02e+03,d_model:512,dropout:0.053,fc_dropout:0.183,patch_size:16,stride:4,seq_len:192
PatchTST CD-192 lr:0.0001,e_layers:3,d_ff:1.02e+03,d_model:128,dropout:0.705,fc_dropout:0.640,patch_size:16,stride:8,seq_len:512
@@
PatchTST CI-336 lr’: 5.207e-05, ’e_layers’: 1, ’d_ff’: 512, ’d_model’: 256, ’dropout’: 0.572, ’fc_dropout’: 0.6308, ’patch_size’: 16, ’stride’: 4, ’seq_len’: 96
PatchTST CD-336 lr’: 2.920e-06, ’e_layers’: 6, ’d_ff’: 1024, ’d_model’: 512, ’dropout’: 0.342, ’fc_dropout’: 0.404, ’patch_size’: 16, ’stride’: 4, ’seq_len’: 192
@@
PatchTST CD-720 lr:0.0024,e_layers:2,d_ff:1.02e+03,d_model:512,dropout:0.450,fc_dropout:0.374,patch_size:16,stride:4,seq_len:512
@@
Table 19. Hyperparameter settings for the ETTh1 dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:3.50e-05,e_layers:1,d_ff:512,d_model:256,dropout:0.571,fc_dropout:0.486,patch_size:16,stride:4,seq_len:96
@@
TSMixer CD-96 lr:0.0011,num_blocks:6,hidden_size:1.02e+03,dropout:0.746,activation:gelu,seq_len:192
@@
PatchTST CI-336 lr’: 3.495e-05, ’e_layers’: 1, ’d_ff’: 512, ’d_model’: 256, ’dropout’: 0.5706, ’fc_dropout’: 0.486, ’patch_size’: 16, ’stride’: 4, ’seq_len’: 96
PatchTST CD-336 lr’: 0.0002, ’e_layers’: 3, ’d_ff’: 512, ’d_model’: 512, ’dropout’: 0.6877, ’fc_dropout’: 0.5936, ’patch_size’: 16, ’stride’: 4, ’seq_len’: 96
@@
TSMixer CI-720 lr:0.0021,num_blocks:6,hidden_size:1.02e+03,dropout:0.141,activation:relu,seq_len:192
@@
Table 20. Hyperparameter settings for the ETTh2 dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:0.0003,e_layers:9,d_ff:256,d_model:128,dropout:0.228,fc_dropout:0.072,patch_size:8,stride:8,seq_len:336
PatchTST CD-96 lr:6.07e-06,e_layers:3,d_ff:256,d_model:1.02e+03,dropout:0.006,fc_dropout:0.642,patch_size:16,stride:8,seq_len:192
TSMixer CI-96 lr:0.0097,num_blocks:2,hidden_size:1.02e+03,dropout:0.115,activation:relu,seq_len:192
@@
PatchTST CI-192 lr:2.97e-06,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.081,fc_dropout:0.435,patch_size:16,stride:8,seq_len:192
PatchTST CD-192 lr:2.97e-06,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.081,fc_dropout:0.435,patch_size:16,stride:8,seq_len:192
TSMixer CI-192 lr:0.0097,num_blocks:2,hidden_size:1.02e+03,dropout:0.115,activation:relu,seq_len:192
@@
PatchTST CI-336 lr’: 0.0003, ’e_layers’: 9, ’d_ff’: 256, ’d_model’: 128, ’dropout’: 0.228, ’fc_dropout’: 0.0716, ’patch_size’: 8, ’stride’: 8, ’seq_len’: 336
PatchTST CD-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.5389, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
TSMixer CI-336 lr:0.0006,num_blocks:8,hidden_size:1.02e+03,dropout:0.166,activation:relu,seq_len:192
@@
PatchTST CI-720 lr:2.27e-06,e_layers:4,d_ff:1.02e+03,d_model:128,dropout:0.198,fc_dropout:0.356,patch_size:16,stride:8,seq_len:512
PatchTST CD-720 lr:0.0019,e_layers:7,d_ff:1.02e+03,d_model:512,dropout:0.471,fc_dropout:0.375,patch_size:8,stride:8,seq_len:720
TSMixer CI-720 lr:0.0097,num_blocks:2,hidden_size:1.02e+03,dropout:0.115,activation:relu,seq_len:192
TSMixer CD-720 lr:0.0011,num_blocks:6,hidden_size:1.02e+03,dropout:0.339,activation:relu,seq_len:192

### v3

@@ Appendix 0.G Full Lookback Window Analysis
Figure 4: Frequency analysis of best-performing lookback windows for forecasting horizon 96.
(a) By Dataset
(b) By CI Models
(c) By CD Models
@@
Figure 5: Frequency analysis of best-performing lookback windows for forecasting horizon 192.
(a) By Dataset
(b) By CI Models
(c) By CD Models
@@
Figure 6: Frequency analysis of best-performing lookback windows for forecasting horizon 336.
(a) By Dataset
(b) By CI Models
(c) By CD Models
@@
Figure 7: Frequency analysis of best-performing lookback windows for forecasting horizon 720.
Appendix 0.H ODE Datasets Sample Visualization
In this section, we provide visualizations for the remaining six ODE datasets utilized in this research. Figure 8 illustrates 1,000 timesteps from the Cell Cycle-based ODE. Specifically, two distinct cyclin degradation activations are observable: channel 2 (middle-top) activates the degradation of cyclin C1 levels (middle-bottom). This phenomenon is replicated on the rightmost two channels for cyclin C2 levels.
Figures 9, 10, 11, and 12 display the first 1,000 timesteps for the Blinking Rotlet, Hopfield, Lorenz, and Lorenz-Coupled datasets, respectively. These visualizations demonstrate how deterministic dynamics generate complex, real-world-like multivariate patterns, where temporal fluctuations in one channel directly influence or trigger corresponding shifts in associated channels.
@@ Appendix 0.H ODE Datasets Sample Visualization
Figure 8: Time series visualization of the cell cycle ODE-based time series sample.
@@
Figure 9: Time series visualization of the blinking rotlet ODE-based time series sample.
Figure 10: Time series visualization of the hopfield ODE-based time series sample.
Figure 11: Time series visualization of the lorenz ODE-based time series sample.
@@
Figure 12: Time series visualization of the Lorenz Coupled ODE-based time series sample.
Appendix 0.I Cell Cycle ODE Discussion
We have already discussed the interesting properties of Double Pendulum ODEs. In this part of the appendix, we would like to further extend the discussion on the coupling between the ODEs and how variables interact across time. Equations 1–6 present the system of ODEs we used to generate the cell cycle dataset [19]. These six equations show two different oscillators based on the concentration levels of cyclin ($C_{1}$ and $C_{2}$) and how they motivate the progression of the cell cycle.
The cyclin concentration reflects directly on the activated Cyclin-Dependent Kinases (cdk1, cdk2), which are represented in the equations by $M_{1}$ and $M_{2}$, respectively; it also directly affects enzymes $X_{1}$ and $X_{2}$. The terms $v_{ij}$ and $v_{dj}$ are constants multiplied in Equation 1 (for $j=1$) and Equation 4 (for $j=2$) as a rate of cyclin synthesis. In Equation 1, we can see clearly the coupling between the ODE variables $C_{1}$ and $M_{2}$, whereas in Equation 4, we see the coupling between $C_{2}$ and $M_{1}$.
All of this demonstrates how information from one channel can cause spikes or changes of trend in other channels, showing clear channel dependence in the generated time-series data. Furthermore, this section is valuable for demonstrating the important application of such equations to a real-life biomedical process, which further elevates the importance of our study. Please note that these equations were taken from [19]; for more details, we encourage the reader to review their work.
$\frac{dC_{1}}{dt}=\frac{v_{i1}K_{im1}}{K_{im1}+M_{2}}-\frac{v_{d1}X_{1}C_{1}}{K_{d1}+C_{1}}-k_{d1}C_{1}$ (7)
$\frac{dM_{1}}{dt}=\frac{V_{1}(1-M_{1})}{K_{1}+(1-M_{1})}-\frac{V_{2}M_{1}}{K_{2}+M_{1}}$ (8)
$\frac{dX_{1}}{dt}=\frac{V_{3}(1-X_{1})}{K_{3}+(1-X_{1})}-\frac{V_{4}X_{1}}{K_{4}+X_{1}}$ (9)
$\frac{dC_{2}}{dt}=\frac{v_{i2}K_{im2}}{K_{im2}+M_{1}}-\frac{v_{d2}X_{2}C_{2}}{K_{d2}+C_{2}}-k_{d2}C_{2}$ (10)
$\frac{dM_{2}}{dt}=\frac{U_{1}(1-M_{2})}{H_{1}+(1-M_{2})}-\frac{U_{2}M_{2}}{H_{2}+M_{2}}$ (11)
$\frac{dX_{2}}{dt}=\frac{U_{3}(1-X_{2})}{H_{3}+(1-X_{2})}-\frac{U_{4}X_{2}}{H_{4}+X_{2}}$ (12)
$V_{1}=\frac{C_{1}}{K_{c1}+C_{1}}V_{M_{1}},\qquad V_{3}=M_{1}\cdot V_{M_{3}}$ (13)
$U_{1}=\frac{C_{2}}{K_{c2}+C_{2}}U_{M_{1}},\qquad U_{3}=M_{2}\cdot U_{M_{3}}$ (14)
Appendix 0.J Ablation Study on RevIN
@@
In this section, we present an ablation study evaluating the impact of Reversible Instance Normalization (RevIN) [11] on two standard datasets and two datasets from the chaotic benchmark. Since Crossformer was originally proposed without RevIN, we integrated the RevIN component into both the Channel-Independent (CI) and Channel-Dependent (CD) versions to measure its effect across different dataset types. Additionally, we included PatchTST as a strong baseline to measure the effect of RevIN on its primary CI version.
We observed two distinct patterns in the results, as detailed in Table 13:
@@
  • The introduction of RevIN improves the performance of both Crossformer variants on the standard datasets.
  • Conversely, on the chaotic ODE datasets, the models without RevIN outperform their counterparts that include the normalization technique.
@@
These findings confirm that, as established in the literature [3], RevIN generally enhances model performance on standard datasets. While this was not universally true for PatchTST in our tests, it holds for most models in existing studies. Importantly, as emphasized throughout this paper, generalizing across different types of datasets is a non-trivial task; RevIN is no exception, as it deteriorates performance when applied to Crossformer on chaotic ODE datasets. This suggests that the complex, non-stationary patterns inherent in chaotic ODEs may not benefit from the normalization-denormalization technique introduced by RevIN.
Table 13: Ablation study on Reversible Instance Normalization (RevIN). The results demonstrate that while RevIN improves performance on standard datasets (ETTh2, ETTm1), it can be detrimental to chaotic ODE datasets (Hopfield, LorenzCoupled). The best variant for each model is in bold, and the best overall result for each dataset is highlighted in blue. Forecasting horizon $L=96$.
Dataset / Model PatchTST (CI) CrossFormer (CD) CrossFormer (CI)
+RevIN –RevIN +RevIN –RevIN +RevIN –RevIN
@@
Hyperparameter settings for the BlinkingRotlet dataset across different forecasting horizons. The table is split across pages for readability.
Model Setting Best Hyperparameters
[…]
Model Setting Best Hyperparameters
[…]
PatchTST CD-96 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
@@
PatchTST CI-192 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
PatchTST CD-192 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
TSMixer CI-192 lr: 0.0005, num_blocks: 8, hidden_size: 1024, dropout: 0.705, activation: relu, seq_len: 512
@@
PatchTST CI-336 lr: 1.96e-07, e_layers: 5, d_ff: 256, d_model: 1024, dropout: 0.295, fc_dropout: 0.622, patch_size: 8, stride: 8, seq_len: 720
PatchTST CD-336 lr: 9.16e-05, e_layers: 6, d_ff: 1024, d_model: 1024, dropout: 0.086, fc_dropout: 0.509, patch_size: 16, stride: 8, seq_len: 192
@@
PatchTST CI-720 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
@@
TSMixer CI-720 lr: 0.0004, num_blocks: 8, hidden_size: 1024, dropout: 0.754, activation: relu, seq_len: 512
@@
Hyperparameter settings for the CellCycle dataset. The table is split across pages for readability.
Model Setting Best Hyperparameters
[…]
Model Setting Best Hyperparameters
[…]
@@
PatchTST CI-336 lr: 0.00027, e_layers: 10, d_ff: 256, d_model: 128, dropout: 0.081, fc_dropout: 0.390, patch_size: 8, stride: 8, seq_len: 336
PatchTST CD-336 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
@@
PatchTST CI-720 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
PatchTST CD-720 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
@@
Hyperparameter settings for the DoublePendulum dataset. The table is split across pages for readability.
Model Setting Best Hyperparameters
[…]
Model Setting Best Hyperparameters
[…]
@@
PatchTST CI-192 lr: 1.53e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.112, fc_dropout: 0.457, patch_size: 16, stride: 4, seq_len: 192
PatchTST CD-192 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
@@
TSMixer CD-192 lr: 0.0032, num_blocks: 7, hidden_size: 1024, dropout: 0.006, activation: relu, seq_len: 192
@@
PatchTST CI-336 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
PatchTST CD-336 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
TSMixer CI-336 lr: 0.0012, num_blocks: 8, hidden_size: 1024, dropout: 0.307, activation: relu, seq_len: 192
@@
PatchTST CD-720 lr: 0.0002, e_layers: 1, d_ff: 1024, d_model: 512, dropout: 0.471, fc_dropout: 0.710, patch_size: 8, stride: 8, seq_len: 336
TSMixer CI-720 lr: 0.0005, num_blocks: 7, hidden_size: 1024, dropout: 0.169, activation: relu, seq_len: 192
@@
Hyperparameter settings for the ETTh1 dataset across different forecasting horizons.
Model Setting Best Hyperparameters
[…]
Model Setting Best Hyperparameters
[…]
@@
PatchTST CI-192 lr: 7.27e-06, e_layers: 4, d_ff: 1024, d_model: 512, dropout: 0.053, fc_dropout: 0.183, patch_size: 16, stride: 4, seq_len: 192
PatchTST CD-192 lr: 0.0001, e_layers: 3, d_ff: 1024, d_model: 128, dropout: 0.705, fc_dropout: 0.640, patch_size: 16, stride: 8, seq_len: 512
@@
PatchTST CI-336 lr: 5.21e-05, e_layers: 1, d_ff: 512, d_model: 256, dropout: 0.572, fc_dropout: 0.631, patch_size: 16, stride: 4, seq_len: 96
PatchTST CD-336 lr: 2.92e-06, e_layers: 6, d_ff: 1024, d_model: 512, dropout: 0.342, fc_dropout: 0.404, patch_size: 16, stride: 4, seq_len: 192
@@
PatchTST CD-720 lr: 0.0024, e_layers: 2, d_ff: 1024, d_model: 512, dropout: 0.450, fc_dropout: 0.374, patch_size: 16, stride: 4, seq_len: 512
@@
Hyperparameter settings for the ETTh2 dataset across different forecasting horizons.
Model Setting Best Hyperparameters
[…]
Model Setting Best Hyperparameters
[…]
@@
TSMixer CD-96 lr: 0.0011, num_blocks: 6, hidden_size: 1024, dropout: 0.746, activation: gelu, seq_len: 192
@@
PatchTST CI-336 lr: 3.50e-05, e_layers: 1, d_ff: 512, d_model: 256, dropout: 0.571, fc_dropout: 0.486, patch_size: 16, stride: 4, seq_len: 96
PatchTST CD-336 lr: 0.0002, e_layers: 3, d_ff: 512, d_model: 512, dropout: 0.688, fc_dropout: 0.594, patch_size: 16, stride: 4, seq_len: 96
@@
TSMixer CI-720 lr: 0.0021, num_blocks: 6, hidden_size: 1024, dropout: 0.141, activation: relu, seq_len: 192
@@
Hyperparameter settings for the ETTm1 dataset across different forecasting horizons.
Model Setting Best Hyperparameters
[…]
Model Setting Best Hyperparameters
[…]
PatchTST CD-96 lr: 6.07e-06, e_layers: 3, d_ff: 256, d_model: 1024, dropout: 0.006, fc_dropout: 0.642, patch_size: 16, stride: 8, seq_len: 192
TSMixer CI-96 lr: 0.0097, num_blocks: 2, hidden_size: 1024, dropout: 0.115, activation: relu, seq_len: 192
@@
PatchTST CI-192 lr: 2.97e-06, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.081, fc_dropout: 0.435, patch_size: 16, stride: 8, seq_len: 192
PatchTST CD-192 lr: 2.97e-06, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.081, fc_dropout: 0.435, patch_size: 16, stride: 8, seq_len: 192
TSMixer CI-192 lr: 0.0097, num_blocks: 2, hidden_size: 1024, dropout: 0.115, activation: relu, seq_len: 192
@@
PatchTST CI-336 lr: 0.0003, e_layers: 9, d_ff: 256, d_model: 128, dropout: 0.228, fc_dropout: 0.0716, patch_size: 8, stride: 8, seq_len: 336
PatchTST CD-336 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.0137, fc_dropout: 0.5389, patch_size: 16, stride: 8, seq_len: 192
TSMixer CI-336 lr: 0.0006, num_blocks: 8, hidden_size: 1024, dropout: 0.166, activation: relu, seq_len: 192
@@
PatchTST CI-720 lr: 2.27e-06, e_layers: 4, d_ff: 1024, d_model: 128, dropout: 0.198, fc_dropout: 0.356, patch_size: 16, stride: 8, seq_len: 512
PatchTST CD-720 lr: 0.0019, e_layers: 7, d_ff: 1024, d_model: 512, dropout: 0.471, fc_dropout: 0.375, patch_size: 8, stride: 8, seq_len: 720
TSMixer CI-720 lr: 0.0097, num_blocks: 2, hidden_size: 1024, dropout: 0.115, activation: relu, seq_len: 192
TSMixer CD-720 lr: 0.0011, num_blocks: 6, hidden_size: 1024, dropout: 0.339, activation: relu, seq_len: 192